HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

VUV PHOTOIONIZATION SPECTROSCOPY OF INTERSTELLAR NITRILES

H.R. HRODMARSSON, \underline{M. SCHWELL} , N. FRAY, Laboratoire Interuniversitaire des Systèmes Atmosphériques, CNRS UMR 7583, Université Paris-Est Créteil et Université Paris Cité, IPSL, 61 av. Général de Gaulle, 94010

Créteil, France; E.-L. ZINS, De la Molécule aux Nano-objets: Réactivité, Interactions et Spectroscopies, CNRS UMR 8233, Sorbonne Université, 4 place Jussieu Tour 53, 75252 Paris cedex 05; J.-C. GUILLEMIN, Institut des Sciences Chimiques de Rennes, CNRS UMR 6226, Université Rennes 1, 263 Avenue Général

Leclerc, 35700 Rennes, France

Our group has been studying the VUV photoionization of interstellar molecules for a decade, focusing on so-called interstellar complex organic molecules (iCOMs) which usually contain more than 5 atoms. Here we study the photoionization dynamics and the dissociative photoionization, via mass-selected threshold photoelectron spectroscopy (TPES) and velocity map imaging (VMI). To derive photoionization and dissociative photoionization rates for astrophysical applications, ionization and dissociative photoionization cross sections can be estimated to better than an order of magnitude. Here, we used the DELICIOUS 3 spectrometer installed as a permanent end station on the DESIRS beamline at the Synchrotron SOLEIL facility in France, to study a series of interstellar nitriles; all of which have been observed in molecular clouds. These are the following: 1) methyl cyano acetylene (MCA; C_{4}H_{3}N), 2) Vinyl cyano acetylene (VCA; C_{5}H_{3}N) and 3) 2-cyano propene (2CP; C_{4}H_{5}N). MCA has been detected since 1984 by several groups ^{1}, whereas VCA and 2CP have been detected only in 2021^{2}^{,}^{3}. Detections are all in the Taurus Molecular Cloud (TMC-1). We will present preliminary results on photoionization mass spectra and slow photoelectron spectra of parent and fragment ions, from the respective ionization energies up to 15.5 eV. This energy range suffices to derive accurate photo rates in cold molecular clouds due to cosmic ray ionization. Calculations are performed too in order to rationalize the chemical structure of the formed fragments. Astrophysical implications of our study will be discussed. This work was supported by the Programme National de Physique et Chimie du Milieu Interstellaire (PCMI) of CNRS/INSU with INC/INP co-funded by CEA and CNES.

1N.W. Broten, J.M. MacLeod, L.W. Avery, W.M. Irvine, B. Hoglund, P. Friberg, and A. Hjalmarson, Astrophys. J. 1984, 276, L25-L29.

2K.L.K. Lee, R.A. Loomis, A.M. Burkhardt, I.R. Cooke, C. Xue, M.A. Siebert, C.N. Shingledecker, A. Remijan, S.B. Charnley, M.C. McCarthy, B.A. McGuire, Astrophys. J. Lett. 2021, 908, L11.

3J. Cernicharo, R. Fuentetaja, C. Cabezas, M. Agundez, N. Marcelino, B. Tercero, J. R. Pardo, P. de Vicente, Astron. Astrophys. 2022, 663, L5.

p-number: p061  Submitted on Fri, 09 Jun 2023 17:05:00 +0200